OpenStudy (shelby1290):

The cost, in dollars, of operating a machine per day is given by the formula C=2t^2-84t+1025, where t is the time, in hours, the machine operates. What is the minimum cost of running the machine? For how many hours must the machine run to reach this minimum cost?

OpenStudy (sepeario):

The minimum time is where t=1 because 1 hour is the least time, hence substitute this value into the polynomial. \[C=2*1-84*1+1025=943\] I'm pretty sure that's the answer.

OpenStudy (sepeario):

If they mean working for a full day, then you will have to substitute 24 hours (a day) into the equation; hence: \[C=2(24*24)-(84*24)+1025=1152-2016+1025=161\] that is probably the minimum because it is less than the original one.

OpenStudy (shelby1290):

@Sepeario at the back of my textbook it says the minimum cost of $143 occurs when the machine runs for 21 hours

OpenStudy (shelby1290):

so...i really don't know what to write

OpenStudy (anonymous):

\[c=2t^2-84t+1025\] \[\frac{ dc }{ dt }=4t-84\] \[\frac{ dc }{ dt }=0\] gives 4t-84=0 t=21 \[\frac{ d^2c }{ dt^2 }=4\] at t=21 \[\frac{ d^2c }{ dt^2 }=4>0\] c is minimum at t=21 so \[minimum~ c=2(21)^2-84*21+1025=?\]

OpenStudy (anonymous):

Or you can also solve by completing the squares.